I have an isoceles with an angle supplementary to angle 3 equaling 116 degrees find angles 1 and 2.
Answer by mananth(16946)

I have an isoceles with an angle supplementary to angle 3 equaling 116 degrees find angles 1 and 2.
supplement to 116 = 64
let the equal angles be x
sum of these two will be 2x
2x+64=180 ( sum of agles of triangle)
2x=180-64
2x=116
x=58
the angles are 58,58 and 64